A light beam containing red and violet wavelengths is incident on a slab of quartz at an angle of incidence of 78.7°. The index of refraction of quartz is 1.455 at 600 nm (red light), and its index of refraction is 1.468 at 410 nm (violet light). Find the dispersion of the slab, which is defined as the difference in the angles of refraction for the two wavelengths.
Answer #1

Solution) Angle of incidence i=78.7°

Index of refraction n=1.455 at 600nm(red)

n=1.468 at 410nm(violet)

Angle of refraction r=?

We have n=sin(i)/sin(r)

For red n=1.455

sin(r)=sin(78.7)/(1.455)=0.673

r=sin inverse(0.673)=42.29°

For violet n'=1.468

Sin(r')=sin(78.7)/1.468=0.667

r'=41.83°

The dispersion =r-r'

Dispersion=42.29-41.83

Dispersion=0.46°